five numbers are formed with 0,1,2,3,4,(not allowing a digit being replaced in any number)the probability of getting 2 in the tens place and 0 in the units place always is

Grade:12

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hiii reddy !

 

Condition is     _ _ _ 2 0 

 

Now we have  1   3   4  ....

The number of ways of creating a no. by using this three digits..without replacement  is   3X2X1 = 6

 

Total no of ways ob obtaining a 5-digit no ..using 0 1 2 3 4 not replacing the no.s   =  4 X 4 X 3 X 2 X 1   = 96

 

Now probabilty =  6/96  = 1/16   ans

 

With regards

 

Yagya

askiitians_expert